OpenStudy (anonymous):

How do you slove 4x+5=21

OpenStudy (anonymous):

-5 both sides, 21-5 = 16 then divide each sife by 4, 16/4. now x =4 check. 4(4) + 5 = 21 16+5 is 21. 4 is answer

OpenStudy (anonymous):

*side